OpenStudy (anonymous):

Two rooms in a house are similar rectangles. Room A is 12 ft by 16 ft. The longer side of Room B is 4 ft shorter than twice the length of the shorter side of Room A. What are the dimensions of the second Room B?

OpenStudy (anonymous):

Area of room A = 12*16 = 192 square foot. Dimensions of Room B: Length = 2*Width - 4 Area of Room B = Width * (2*Width-1) = 192 Solve.

OpenStudy (anonymous):

So what would the answer be? would it be 16 * 16 14 * 16 16 * 20 15* 20

OpenStudy (anonymous):

You need to solve that equation about to find the width and then the length.

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

Solve the following for x:\[\frac{12}{16}=\frac{x}{2*12-4}\]x = 15, room is 15 by 20 ft